# Random Lévy Looptrees and Lévy Maps

### v2

@@ 3.5.   Proof of the technical lemma
$-r<W_{G_{k}}\leq r\log^{10k+3}(1/r)\qquad\text{and so}\qquad\Delta_{k}\leq 3(r\log^{10k+6}(1/r))^{2}\qquad\text{for each}\enskip k\leq K_{r}.$
According to the strong Markov property applied at the first hitting time of $-r\log^{10k}(1/r)$, the path $W$ behaves after this time as an unconditioned Brownian motion started from this value during the remaining time of the interval $[G_{k},D_{k}]$, which is at most $3(r\log^{10k+6}(1/r))^{2}$ on the previous event. This is not enough for $W$ to further reach $-r\log^{10(k+1)}(1/r)$. Indeed, if $X$ is a generic Brownian motion, then:
$\mathop{}\!\mathbb{P}\Bigl{(}\inf_{[0,4(r\log^{10k+6}(1/r))^{2}]}X<-r\log^{10k+8}(1/r)\Bigr{)}=\mathop{}\!\mathbb{P}\Bigl{(}|X_{1}|>\frac{1}{2}\log^{2}(1/r)\Bigr{)}\leq 2\exp\Bigl{(}-\frac{1}{4}\log^{2}(1/r)\Bigr{)},$
@@
Step 2: Proof of for Brownian bridges. Let us next prove the actual statement by comparing the Brownian motions $X^{L,k}$ and $X^{R,k}$ with Brownian bridges $W^{L,k}$ and $W^{R,k}$. We aim at proving that the probability that one of these bridges stays above $-r$ in the first half of the interval when $\Delta_{k}>(r+W_{G_{k}})^{2}\log^{6}(1/r)$ is summable when $r$ ranges over $\{2^{-n},n\in\mathbb{N}\}$, and then that if both $W_{G_{k}}\leq r\log^{10k}(1/r)$ and $\Delta_{k}>(r\log^{10k}(1/r))^{2}$, then there is a probability bounded away from 00 uniformly in $r$ and $k$ that the two bridges both reach $-r$. Let us only consider $X^{L,k}$ and $W^{L,k}$ since the same argument applies to $X^{R,k}$ and $W^{R,k}$. Since, conditionally on $W_{G_{k}}$, $W_{D_{k}}$, $\Delta_{k}$, and $\Delta^{L}_{k}$, the path $W^{L,k}$ is a Brownian bridge from $W_{G_{k}}$ to $W_{D_{k}}$ with duration $\Delta^{L}_{k}\geq\Delta_{k}$, then the following absolute continuity relation holds: for any measurable and nonnegative function $F$,
$\mathop{}\!\mathbb{E}\bigl{[}F(W^{L,k}_{t},t\leq\Delta^{L}_{k}/2)\mid W_{G_{k}},W_{D_{k}},\Delta_{k},\Delta^{L}_{k}\bigr{]}=\mathop{}\!\mathbb{E}\Bigl{[}F(X^{L,k}_{t},t\leq\Delta^{L}_{k}/2)\,\frac{p_{\Delta^{L}_{k}/2}(W_{D_{k}}-X^{L}_{\Delta^{L}_{k}/2})}{p_{\Delta^{L}_{k}}(W_{D_{k}}-W_{G_{k}})}\bigm{|}W_{G_{k}},W_{D_{k}},\Delta_{k},\Delta^{L}_{k}\Bigr{]},$
@@
This allows to infer from the first part of the proof of that
$\mathop{}\!\mathbb{P}\Bigl{(}\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)\enskip\text{and}\enskip\max\Bigl{\{}\inf_{t\in[0,\Delta_{k}/2]}W^{L,k}_{t},\inf_{t\in[0,\Delta_{k}/2]}W^{R,k}_{t}\Bigr{\}}>-r\Bigr{)}=O(\log^{-3}(1/r)),$
and then that, as in the first part, we must impose $\Delta_{k}\leq 2(r+W_{G_{k}})^{2}\log^{6}(1/r)\leq 3(r\log^{10k+6}(1/r))^{2}$ for every $k\leq K_{r}$ in order to prevent the bridges $W^{L,k}$ and $W^{R,k}$ to reach $-r$.
Let us prove the upper bound (3.8). Conditionally on $W_{G_{k}}$ and $\Delta_{k}$, the path $(W_{t};t\in[G_{k},D_{k}])$ has the law of a Brownian motion $Y$ with duration $\Delta_{k}$, started from $W_{G_{k}}$, and conditioned on the event $E_{k}=\{\inf\{Y_{t};t\in[G_{k},D_{k}]\}\leq-r\log^{10k}(1/r)\}$. Then the law of $W_{D_{k}}$ can be obtained by splitting at the first hitting time of this value; by the Markov, property, the remaining path after is an unconditioned Brownian motion. Recall also that we work implicitly on the event where $W_{G_{k}},W_{D_{k}}>-r$. For $z>0$, let us write $q_{z}$ for the density of the first hitting time of $-z$ by a standard Brownian motion, then:
@@
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}t\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ q_{r\log^{10k}(1/r)+W_{G_{k}}}(t)p_{\Delta_{k}-t}(y+r\log^{10k}(1/r))\frac{p_{\Delta^{L}_{k}/2}(y-X^{L}_{\Delta^{L}_{k}/2})}{p_{\Delta^{L}_{k}}(y-W_{G_{k}})}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}t\frac{q_{r\log^{10k}(1/r)+W_{G_{k}}}(t)}{2\sqrt{\pi(\Delta_{k}-t)}}\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r\log^{10k}(1/r))^{2}}{2(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}.$
We have $\mathop{}\!\mathbb{P}(E_{k})=\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}tq_{r\log^{10k}(1/r)+W_{G_{k}}}(t)$, it thus remains to prove that the integral in $y$ of the rest of the expression is bounded, when in addition $\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)$. Let us first consider the integral over $[W_{G_{k}},\infty)$. Indeed, if $y>W_{G_{k}}>-r$, then $(y-W_{G_{k}})^{2}\leq y^{2}\leq(y+r\log^{10k}(1/r))^{2}$ and consequently
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r\log^{10k}(1/r))^{2}}{2(\Delta_{k}-t)}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}.$
@@
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{4(\Delta_{k}-t)}\Bigr{)},$
@@
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2}\frac{t}{\Delta_{k}(\Delta_{k}-t)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{2\Delta_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi\Delta_{k}/2}}\exp\Bigl{(}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{2\Delta_{k}}\Bigr{)}\mathbf{1}_{\{0<t\leq\Delta_{k}/2\}}+\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{4(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{\Delta_{k}/2<t<\Delta_{k}\}}.$
Again the integral over the entire real line is constant. Finally, if $-r<y<W_{G_{k}}$ then $(y-W_{G_{k}})^{2}\leq(r+W_{G_{k}})^{2}$. Here we use that we work on the event $\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)\geq 2(y-W_{G_{k}})^{2}\log^{6}(1/r)$. Indeed, this implies:
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r\log^{10k}(1/r))^{2}}{2(\Delta_{k}-t)}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{1}{4\log^{6}(1/r)}-\frac{(y+r\log^{10k}(1/r))^{2}}{2(\Delta_{k}-t)}\Bigr{)},$
@@
Let us finally lower bound the probability that the Brownian bridge $W^{L,k}$ reaches $-r$ in the first half of the interval when both $-r<W_{G_{k}}\leq r\log^{10k}(1/r)$ and $\Delta_{k}>(r\log^{10k}(1/r))^{2}$. We argue similarly: conditionally on $\Delta^{L}_{k}$, $W_{G_{k}}$, and $W_{D_{k}}$, this probability equals
$\int_{0}^{\Delta^{L}_{k}/2}\mathop{}\!\mathrm{d}t\ q_{-(r+W_{G_{k}})}(t)\frac{p_{\Delta^{L}_{k}-t}(W_{D_{k}}+r)}{p_{\Delta^{L}_{k}}(W_{D_{k}}-W_{G_{k}})}$ $=\int_{0}^{\Delta^{L}_{k}/2}\mathop{}\!\mathrm{d}t\ q_{-(r+W_{G_{k}})}(t)\sqrt{\frac{\Delta^{L}_{k}}{\Delta^{L}_{k}-t}}\exp\Bigl{(}\frac{(W_{D_{k}}-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(W_{D_{k}}+r)^{2}}{2(\Delta^{L}_{k}-t)}\Bigr{)}$
@@
Without the exponential, the last integral is bounded away from 00. We then lower bound the expectation of the exponential with respect to $W_{D_{k}}$. Recall that conditionally on $W_{G_{k}}$ and $\Delta_{k}$, the path $(W_{t};t\in[G_{k},D_{k}])$ has the law of a Brownian motion started from $W_{G_{k}}$ and conditioned on the event $E_{k}$ that it reaches $-r\log^{10k}(1/r)$ at some point in the time interval, so by splitting at the first hitting time of this value, we have since we restrict our attention to $t\leq\Delta_{k}/2$,
@@
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r\log^{10k}(1/r)+W_{G_{k}})}(s)}{\sqrt{2\pi(\Delta_{k}-s)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2(\Delta_{k}-s)}-\frac{(y+r)^{2}}{2(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$\geq\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r\log^{10k}(1/r)+W_{G_{k}})}(s)}{\sqrt{2\pi(\Delta_{k}-s)}}\exp\Bigl{(}-\frac{(y+r\log^{10k}(1/r))^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-s}+\frac{1}{\Delta_{k}-t}\Bigr{)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$\geq\int_{0}^{\Delta_{k}/2}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r\log^{10k}(1/r)+W_{G_{k}})}(s)}{\sqrt{2\pi\Delta_{k}}}\exp\Bigl{(}-2\frac{(y+r\log^{10k}(1/r))^{2}}{\Delta_{k}}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}.$
Consider the integral over $y$ only in the last line, and without de density $q$. It equals half the probability that a Gaussian random variable with mean $-r\log^{10k}(1/r)$ and variance $\Delta_{k}/4$ is larger than $-r$, or equivalently that a Gaussian random variable with mean 00 and variance $1$ is larger than $2(r\log^{10k}(1/r)-r)/\sqrt{\Delta_{k}}$. Since $\Delta_{k}>(r\log^{10k}(1/r))^{2}$, this probability is indeed bounded away from 00 uniformly in $k$ and $r$. Then the integral of $q$ over $s$ equals the probability that a standard Brownian motion reaches $-(W_{G_{k}}+r\log^{10k}(1/r))$ in less than $\Delta_{k}/2$ when $W_{G_{k}}\leq r\log^{10k}(1/r)\leq\sqrt{\Delta_{k}}$, which is again bounded away from 00.
Combined with the results on $X^{L,k}$, we infer that there exists some constant $c>0$ such that for all $r<1$, on each interval $[G_{k},D_{k}]$ with $W_{G_{k}}\leq r\log^{10k}(1/r)$ and $\Delta_{k}\geq(r\log^{10k}(1/r))^{2}$, there is a probability at least $c$ that $W^{L,k}$ reaches the value $-r$ in the first half of the interval. The same holds for $X^{L,k}$ and since they are independent, they both reach the value $-r$ in the first half with probability at least $c^{2}$. We can then conclude from Lemma 3.12 that this occurs for a positive proportion of indices $k\leq K_{r}$. ∎

### v3

@@ 4.5.   Proof of the technical lemma
$-r<W_{G_{k}}\leq r\log^{10k+3}(1/r)\quad\text{and so}\quad\Delta_{k}\leq 3(r\log^{10k+6}(1/r))^{2}\quad\text{for each}\enskip k\leq K_{r}.$
According to the strong Markov property applied at the first hitting time of $-r\log^{10k}(1/r)$, the path $W$ behaves after this time as an unconditioned Brownian motion started from this value during the remaining time of the interval $[G_{k},D_{k}]$. On the previous event, this remaining time is at most $3(r\log^{10k+6}(1/r))^{2}$. This is not enough for $W$ to further reach $-r\log^{10(k+1)}(1/r)$. Indeed, if $X$ is a generic Brownian motion, then:
$\mathop{}\!\mathbb{P}\Bigl{(}\inf_{[0,4(r\log^{10k+6}(1/r))^{2}]}X<-r\log^{10k+8}(1/r)\Bigr{)}$ $=\mathop{}\!\mathbb{P}\Bigl{(}|X_{1}|>\frac{1}{2}\log^{2}(1/r)\Bigr{)}$
$\leq 2\operatorname{e}^{-\frac{1}{4}\log^{2}(1/r)},$
@@
Step 2: Proof for Brownian bridges. Let us next prove the actual statement by comparing the Brownian motions $X^{L,k}$ and $X^{R,k}$ with Brownian bridges $W^{L,k}$ and $W^{R,k}$. We aim at proving that the probability that one of these bridges stays above $-r$ in the first half of the interval when $\Delta_{k}>(r+W_{G_{k}})^{2}\log^{6}(1/r)$ is summable when $r$ ranges over $\{2^{-n},n\in\mathbb{N}\}$, and then that if both $W_{G_{k}}\leq r\log^{10k}(1/r)$ and $\Delta_{k}>(r\log^{10k}(1/r))^{2}$, then there is a probability bounded away from 00 uniformly in $r$ and $k$ that the two bridges both reach $-r$. Let us only consider $X^{L,k}$ and $W^{L,k}$ since the same argument applies to $X^{R,k}$ and $W^{R,k}$. Since, conditionally on $W_{G_{k}}$, $W_{D_{k}}$, $\Delta_{k}$, and $\Delta^{L}_{k}$, the path $W^{L,k}$ is a Brownian bridge from $W_{G_{k}}$ to $W_{D_{k}}$ with duration $\Delta^{L}_{k}\geq\Delta_{k}$, then the following absolute continuity relation holds: for any measurable and nonnegative function $F$,
$\mathop{}\!\mathbb{E}\bigl{[}F(W^{L,k}_{t},t\leq\Delta^{L}_{k}/2)\mid W_{G_{k}},W_{D_{k}},\Delta_{k},\Delta^{L}_{k}\bigr{]}$
$\qquad=\mathop{}\!\mathbb{E}\Bigl{[}F(X^{L,k}_{t},t\leq\Delta^{L}_{k}/2)\,\frac{p_{\Delta^{L}_{k}/2}(W_{D_{k}}-X^{L}_{\Delta^{L}_{k}/2})}{p_{\Delta^{L}_{k}}(W_{D_{k}}-W_{G_{k}})}\bigm{|}W_{G_{k}},W_{D_{k}},\Delta_{k},\Delta^{L}_{k}\Bigr{]},$
@@
This allows to infer from the first part of the proof that
$\mathop{}\!\mathbb{P}\Bigl{(}\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)\enskip\text{and}\enskip\max\Bigl{\{}\inf_{t\in[0,\Delta_{k}/2]}W^{L,k}_{t},\inf_{t\in[0,\Delta_{k}/2]}W^{R,k}_{t}\Bigr{\}}>-r\Bigr{)}$
is $O(\log^{-3}(1/r))$. Then, as in the first part, we must impose $\Delta_{k}\leq 2(r+W_{G_{k}})^{2}\log^{6}(1/r)\leq 3(r\log^{10k+6}(1/r))^{2}$ for every $k\leq K_{r}$ in order to prevent the bridges $W^{L,k}$ and $W^{R,k}$ to reach $-r$.
Let us prove the upper bound (4.8). Recall the notation $r_{k}=r\log^{10k}(1/r)$. Conditionally on $W_{G_{k}}$ and $\Delta_{k}$, the path $(W_{t};t\in[G_{k},D_{k}])$ has the law of a Brownian motion $Y$ with duration $\Delta_{k}$, started from $W_{G_{k}}$, and conditioned on the event $E_{k}=\{\inf\{Y_{t};t\in[G_{k},D_{k}]\}\leq-r_{k}\}$. Then the law of $W_{D_{k}}$ can be obtained by splitting at the first hitting time of this value; by the Markov, property, the remaining path after is an unconditioned Brownian motion. Recall also that we work implicitly on the event where $W_{G_{k}},W_{D_{k}}>-r$. For $z>0$, let us write $q_{z}$ for the density of the first hitting time of $-z$ by a standard Brownian motion, then:
@@
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}t\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ q_{r_{k}+W_{G_{k}}}(t)p_{\Delta_{k}-t}(y+r_{k})\frac{p_{\Delta^{L}_{k}/2}(y-X^{L}_{\Delta^{L}_{k}/2})}{p_{\Delta^{L}_{k}}(y-W_{G_{k}})}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}t\frac{q_{r_{k}+W_{G_{k}}}(t)}{2\sqrt{\pi(\Delta_{k}-t)}}$
$\qquad\qquad\qquad\cdot\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r_{k})^{2}}{2(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}.$
We have $\mathop{}\!\mathbb{P}(E_{k})=\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}tq_{r_{k}+W_{G_{k}}}(t)$, it thus remains to prove that the integral in $y$ of the rest of the expression is bounded, when in addition $\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)$. Let us first consider the integral over $[W_{G_{k}},\infty)$. Indeed, if $y>W_{G_{k}}>-r$, then $(y-W_{G_{k}})^{2}\leq y^{2}\leq(y+r_{k})^{2}$ and consequently
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r_{k})^{2}}{2(\Delta_{k}-t)}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}.$
@@
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{4(\Delta_{k}-t)}\Bigr{)},$
@@
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-t}-\frac{1}{\Delta^{L}_{k}}\Bigr{)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2}\frac{t}{\Delta_{k}(\Delta_{k}-t)}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{2\Delta_{k}}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi\Delta_{k}/2}}\exp\Bigl{(}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{2\Delta_{k}}\Bigr{)}\mathbf{1}_{\{0<t\leq\Delta_{k}/2\}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{4(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{\Delta_{k}/2<t<\Delta_{k}\}}.$
Again the integral over the entire real line of the upper bound is constant. Finally, if $-r<y<W_{G_{k}}$ then $(y-W_{G_{k}})^{2}\leq(r+W_{G_{k}})^{2}$. Here we use that we work on the event $\Delta_{k}>2(r+W_{G_{k}})^{2}\log^{6}(1/r)\geq 2(y-W_{G_{k}})^{2}\log^{6}(1/r)$. Indeed, this implies:
$\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{(y-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(y-X^{L}_{\Delta^{L}_{k}/2})^{2}}{\Delta^{L}_{k}}-\frac{(y+r_{k})^{2}}{2(\Delta_{k}-t)}\Bigr{)}$
$\leq\frac{1}{2\sqrt{\pi(\Delta_{k}-t)}}\exp\Bigl{(}\frac{1}{4\log^{6}(1/r)}-\frac{(y+r_{k})^{2}}{2(\Delta_{k}-t)}\Bigr{)},$
@@
Let us finally lower bound the probability that the Brownian bridge $W^{L,k}$ reaches $-r$ in the first half of the interval when both $-r<W_{G_{k}}\leq r_{k}$ and $\Delta_{k}>r_{k}^{2}$, where we keep the notation $r_{k}=r\log^{10k}(1/r)$. We argue similarly: conditionally on $\Delta^{L}_{k}$, $W_{G_{k}}$, and $W_{D_{k}}$, this probability equals
$\int_{0}^{\Delta^{L}_{k}/2}\mathop{}\!\mathrm{d}t\ q_{-(r+W_{G_{k}})}(t)\frac{p_{\Delta^{L}_{k}-t}(W_{D_{k}}+r)}{p_{\Delta^{L}_{k}}(W_{D_{k}}-W_{G_{k}})}$
$=\int_{0}^{\Delta^{L}_{k}/2}\mathop{}\!\mathrm{d}t\ q_{-(r+W_{G_{k}})}(t)\sqrt{\frac{\Delta^{L}_{k}}{\Delta^{L}_{k}-t}}\exp\Bigl{(}\frac{(W_{D_{k}}-W_{G_{k}})^{2}}{2\Delta^{L}_{k}}-\frac{(W_{D_{k}}+r)^{2}}{2(\Delta^{L}_{k}-t)}\Bigr{)}$
@@
Without the exponential, the last integral is bounded away from 00. We then lower bound the expectation of the exponential with respect to $W_{D_{k}}$. Recall that conditionally on $W_{G_{k}}$ and $\Delta_{k}$, the path $(W_{t};t\in[G_{k},D_{k}])$ has the law of a Brownian motion started from $W_{G_{k}}$ and conditioned on the event $E_{k}$ that it reaches $-r_{k}$ at some point in the time interval, so by splitting at the first hitting time of this value, we have since we restrict our attention to $t\leq\Delta_{k}/2$,
@@
$=\frac{1}{\mathop{}\!\mathbb{P}(E_{k})}\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r_{k}+W_{G_{k}})}(s)}{\sqrt{2\pi(\Delta_{k}-s)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2(\Delta_{k}-s)}-\frac{(y+r)^{2}}{2(\Delta_{k}-t)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$\geq\int_{0}^{\Delta_{k}}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r_{k}+W_{G_{k}})}(s)}{\sqrt{2\pi(\Delta_{k}-s)}}\exp\Bigl{(}-\frac{(y+r_{k})^{2}}{2}\Bigl{(}\frac{1}{\Delta_{k}-s}+\frac{1}{\Delta_{k}-t}\Bigr{)}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}$
$\geq\int_{0}^{\Delta_{k}/2}\mathop{}\!\mathrm{d}s\int_{-r}^{\infty}\mathop{}\!\mathrm{d}y\ \frac{q_{-(r_{k}+W_{G_{k}})}(s)}{\sqrt{2\pi\Delta_{k}}}\exp\Bigl{(}-2\frac{(y+r_{k})^{2}}{\Delta_{k}}\Bigr{)}\mathbf{1}_{\{W_{G_{k}}>-r\}}.$
Consider the integral over $y$ only in the last line, and without de density $q$. It equals half the probability that a Gaussian random variable with mean $-r_{k}$ and variance $\Delta_{k}/4$ is larger than $-r$, or equivalently that a Gaussian random variable with mean 00 and variance $1$ is larger than $2(r_{k}-r)/\sqrt{\Delta_{k}}$. Since $\Delta_{k}>r_{k}^{2}$, this probability is indeed bounded away from 00 uniformly in $k$ and $r$. Then the integral of $q$ over $s$ equals the probability that a standard Brownian motion reaches $-(W_{G_{k}}+r_{k})$ in less than $\Delta_{k}/2$ when $W_{G_{k}}\leq r_{k}\leq\sqrt{\Delta_{k}}$, which is again bounded away from 00.
Combined with the results on $X^{L,k}$, we infer that there exists some constant $c>0$ such that for all $r<1$, on each interval $[G_{k},D_{k}]$ with $W_{G_{k}}\leq r_{k}$ and $\Delta_{k}\geq(r_{k})^{2}$, there is a probability at least $c$ that $W^{L,k}$ reaches the value $-r$ in the first half of the interval. The same holds for $X^{L,k}$ and since they are independent, they both reach the value $-r$ in the first half with probability at least $c^{2}$. We can then conclude from Lemma 4.12 that this occurs for a positive proportion of indices $k\leq K_{r}$. ∎